# Learning Interpretable Hierarchical Dynamical Systems Models from Time Series Data

### v1

@@ 5 Conclusions
While DSR is currently a burgeoning field, the question of how to best integrate dynamically diverse systems into a common structure, and how to harvest this for transfer learning, generalization, and classification, only recently gained traction (Yin et al., 2021; Kirchmeyer et al., 2022; Göring et al., 2024). Unlike TS models, where the interest lies mainly in forecasting or TS classification/ regression, in DSR we aim for generative models that capture invariants of the true, underlying system’s state space. This commonly requires special training algorithms or loss functions (Mikhaeil et al., 2022; Platt et al., 2023). Using GTF (Hess et al., 2023) for training, we demonstrated that hierarchically structured models significantly improve DSR quality and temporal agreement compared to ‘traditional’, non-hierarchical methods. This is a strong indication that our method was able to utilize information from all DS observed, even if located in different dynamical regimes, to boost performance on each individual system. This was confirmed in transfer learning experiments, where models trained on multiple subjects generalized effectively to new conditions with minimal additional data, and did so much more convincingly than various competing methods. Identifying dynamical regimes, instead of just utilizing TS waveform features, is also likely to be a more principled approach to TS foundation models, since it is the underlying dynamics which determine a system’s temporal properties across contexts. This idea is supported by the strong classification results on human EEG data, where subject groups were much more clearly segregated in the DS feature space than in spaces constructed by more ‘conventional’ TS models. These DS feature spaces furthermore turned out to be surprisingly interpretable. Extracting these dynamically most crucial features and grouping subjects along these dimensions may be of considerable interest in areas like medicine or neuroscience, and is a new application that has not been considered in the DSR field so far.
@@
The highly interpretable nature of the subject-level feature spaces, despite the strongly nonlinear and chaotic systems considered, is surprising. In future work, it needs to be determined how general this property is and whether it also holds for other physical and biological systems. It appears that our method automatically infers the critical control parameters of a system that best differentiate between dynamical regimes (by controlling the underlying bifurcation structure), but it is currently unclear what the mathematical basis for this might be. Another potential shortcoming might be that our approach currently is completely unsupervised, while it may be beneficial to use class labels when available. Finally, one useful extension may be a fully Bayesian framework which yields proper posterior uncertainty estimates for parameters.
@@
Code for this paper is available https://anonymous.4open.science/r/hier-shPLRNN-5E3C, and will be made publically available on GitHub with publication at a peer-reviewed journal.
@@ Acknowledgements
This work was funded by the European Union’s Horizon 2020 programme under grant agreement 945263 (IMMERSE), by the Federal Ministry of Science, Education, and Culture (MWK) of the state of Baden-Württemberg within the AI Health Innovation Cluster Initiative and living lab (grant number 31-7547.223-7/3/2), by the German Research Foundation (DFG) within the collaborative research center TRR-265 (project A06 & B08) and Germany’s Excellence Strategy EXC 2181/1 – 390900948 (STRUCTURES), and by the Hector-II foundation.
@@ Appendix A Methodological Details
Figure 7: Illustration of the hierarchization framework.
@@ A.1.1 Generalized Teacher Forcing (GTF)
with $0\leq\alpha\leq 1$, which may be set by line search or adaptively adjusted throughout training. Hess et al. [2023] prove that an optimal choice for $\alpha$ leads to bounded gradients as sequence length $T$ goes to infinity. An extension of GTF to multimodal and non-Gaussian data (multimodal teacher forcing, MTF) has recently been proposed [Brenner et al., 2024b], and applied to short multi-subject non-Gaussian time series [Brenner et al., 2024a]. MTF can be naturally incorporated into our hierarchical framework and hence is an interesting avenue for further applications.
A.1.2 Batching
@@
Given the complexity of the datasets considered here, an optimal batching strategy for stochastic optimization is not obvious a priori. We hence investigated how different batch sizes impact the reconstruction performance of models trained on the 64-subject Lorenz-63 dataset (see main text). The blue curve in Fig. 8 illustrates how performance varies across a range of batch sizes when training with 50 batches per epoch, corresponding to a fixed number of gradient updates per epoch. In this setup, models trained with smaller batch sizes naturally see less data per epoch than those with larger batches. To ensure that the observed effects are not merely due to reduced data exposure with smaller batch sizes, we repeated the experiment using the same batch sizes, but ensuring the entire dataset was seen in each epoch (orange curve). Our findings suggest that smaller batch sizes do not contain enough statistical information across all subjects, leading to higher variance in gradient updates and, in some cases, failure of training. For larger batch sizes, the performance difference was marginal.
@@
The previous experiments were conducted with a fixed number of subjects. However, since we aim to incorporate information from all subjects within each batch, we extended the analysis to examine how the optimal batch size scales with different numbers of subjects. To do this, we randomly sampled smaller subsets from the dataset and repeated the experiment. Fig. 9 gives a heat map of the results, showing that the optimal batch size falls between 4 and 16 times the number of subjects in the dataset. For datasets with a large number of subjects, this range becomes computationally infeasible, so we instead sample batch sequences from a random subset of subjects on each iteration.
A.1.3 Hierarchization scheme
@@
As discussed in Sect. 1, our hierarchization framework is flexible, allowing for different forms of the group-level function $G_{\bm{\theta}_{\text{group}}}$ that maps subject-specific feature vectors to model weights. Here we discuss different choices for parameterizing $G_{\bm{\theta}_{\text{group}}}$ and points to consider in this context. The naive approach, employing direct linear projections, results in over-parameterization. For example, the weight matrices $\bm{W}_{1}$ and $\bm{W}_{2}$ of the shPLRNN have $M\times L$ degrees of freedom, but when parameterized with feature vectors through $\bm{W}_{1}^{(j)}:=\text{mat}(\bm{l}^{(j)}\cdot\bm{P}_{W_{1}},M,L)$, the number of trainable parameters expands to $N_{\text{feat}}\times(M\times L+1)$, which with $N_{\text{feat}}>1$ exceeds the $M\times L$ available entries in $\bm{W}_{1}$.
@@
(analogously for $\bm{W}_{2}$), while the diagonal of $\bm{A}$ and vectors $\bm{b}_{1},\bm{b}_{2}$ were specified directly through linear maps. This reduces the number of parameters to $N_{\text{feat}}\times(M+L+1)$. In the cases where $N_{\text{feat}}<M$, such that the matrices would not be full rank, we applied an extra step to map the feature vector to an auxiliary vector $\mathbb{R}^{M}\ni\bm{a}^{(j)}=\bm{R}\bm{l}^{(j)}$ via the shared [^3] To avoid increasing the number of free parameters, $\bm{R}$ could also be chosen to be random and constant. matrix $\bm{R}\in\mathbb{R}^{M\times N_{\text{feat}}}$, and subsequently mapped to the weights $\bm{W}_{1}^{(j)}=\bm{P}_{W_{1}}\text{diag}(\bm{a}^{(j)})\bm{Q}_{W_{1}}$ (analogously for $\bm{W}_{2}$). We also tested using an MLP for mapping the feature vectors onto the subject-specific DSR model weights, with MLP parameters shared among subjects:
@@
Outer product $3348$ $0.36\pm 0.03$ $0.99$ $16\%$ $0.0986\pm 0.0025$ $-0.63$ $73\%$
MLP $36713$ $0.85\pm 0.09$ $-7.55$ $100\%$ $0.111\pm 0.004$ $-5.54$ $100\%$
A.1.4 Dataset size
To further evaluate the robustness of our approach, we conducted additional experiments on the Lorenz-63 dataset as in Sect. 4.1, but with a varying number of time steps for each subject/system, where $1000$ training time steps corresponds to the setting in Sect. 4.1. Fig. 10 shows the ratio between the hierarchical model’s performance and the baseline performance (with values below $1$ indicating superior performance by the hierarchical model). Even for long training sequences the hierarchical model consistently outperforms the baseline, particularly in terms of state-space divergence.
@@ A.1.4 Dataset size
Figure 10: Relative performance (hierarchical/baseline) as a function of time series length for the 64-subject Lorenz-63 dataset.
A.1.5 Uncertainty Estimates
@@
To obtain the uncertainty estimates (error bars) in Figs.5b and c, we first drew a long trajectory with $10000$ time steps from the ground truth system using ${\rho}^{(S+1)}$. We then randomly sampled multiple training subsequences of length $T_{\text{max}}$, inferred the feature vectors $\hat{\bm{l}}^{(S+1)}$ for each subsequence individually, and calculated the standard deviation of the resulting $\hat{\rho}^{(S+1)}$ estimates. As shown in Fig. 11, the estimates sharpen as longer training sequences are used.
@@
Figure 11: a: Distributions $p(\hat{\rho}^{(j)}\mid\bm{x}_{1...T_{\text{max}}}^{(j)})$ (approximated using kernel density estimates), estimated from $10$ time series each with length with different lengths $T_{\text{max}}$ (color-coded) for the example $\rho^{(j)}_{\text{test}}$ from Fig. 5. b: Corresponding GTF loss curves for different training sequence lengths, illustrating that longer sequences/more training data lead to smaller standard deviation over loss curves.
A.1.6 Hyperparameters

### v2

@@ 5 Conclusions
While DSR is currently a burgeoning field, the question of how to best integrate dynamically diverse systems into a common structure, and how to harvest this for transfer learning, generalization, and classification, only recently gained traction (Yin et al., 2021; Kirchmeyer et al., 2022; Göring et al., 2024). Unlike TS models, where the interest lies mainly in forecasting or TS classification/ regression, in DSR we aim for generative models that capture invariants of the underlying system’s state space. Using GTF (Hess et al., 2023) for training, we demonstrated that hierarchically structured models significantly improve DSR quality compared to non-hierarchical methods. This is a strong indication that our method was able to utilize information from all DS observed, even if located in different dynamical regimes, or if described by different systems of equations, to boost performance on each individual system. This was confirmed in transfer learning experiments, where models trained on multiple subjects generalized effectively to new conditions with minimal additional data, and did so much more convincingly than various competing methods. Identifying dynamical regimes, instead of just utilizing TS waveform features, is also likely to be a more principled approach to TS foundation models, since it is the underlying dynamics which determine a system’s temporal properties across contexts. This idea is supported by the strong classification results on human EEG data, where subject groups were much more clearly segregated in the DS feature space than in spaces constructed by ‘conventional’ TS models. These DS feature spaces further turned out to be surprisingly interpretable. Extracting these dynamically most crucial features and grouping subjects along these dimensions may be of considerable interest in areas like medicine or neuroscience, and is a new application that has not been considered in the DSR field so far.
@@
The highly interpretable nature of the subject-level feature spaces is surprising. It appears that our method automatically infers critical control parameters of a system that best differentiate between dynamical regimes (by controlling the underlying bifurcation structure), but it is currently unclear what the mathematical basis for this might be. Another open issue is how DS could best be integrated into the same model that are related but live on widely different time scales, e.g. brain signals over minutes vs. behavioral monitoring across weeks. While Fig. 28 and the empirical examples confirm our approach provides a means to examine how diverse systems are dynamically related, this currently is limited to DS evolving on similar time scales. In Fig. 16 we illustrated how the hier-shPLRNN could be expanded to capture non-stationarity and drivers in non-autonomous DS in an interpretable way; yet, for specific physical or biological processes this may need to be hooked up with more domain knowledge to establish a mechanistic link to the underlying physics/ biology. A useful extension thus could be the incorporation of prior physical knowledge as in SINDy or PINNs. While the latter is no principle obstacle, as shown here for Markov Neural Operators, the tricky part may be to formulate priors that work across all observed systems and domains.
@@
The code for this paper is publicly available at https://github.com/DurstewitzLab/HierarchicalDSR.
@@ Acknowledgements
This work was funded by the European Union’s Horizon 2020 programme under grant agreement 945263 (IMMERSE), by the German Ministry for Education & Research (BMBF) within the FEDORA (01EQ2403F) consortium, by the Federal Ministry of Science, Education, and Culture (MWK) of the state of Baden-Württemberg within the AI Health Innovation Cluster Initiative and living lab (grant number 31-7547.223-7/3/2), by the German Research Foundation (DFG) within the collaborative research center TRR-265 (project A06 & B08) and by the Hector-II foundation.
@@ Appendix A Methodological Details
Figure 8: Illustration of the training algorithm for one example training sequence.
@@ A.1.1 Generalized Teacher Forcing (GTF)
with $0\leq\alpha\leq 1$, which may be set by line search or adaptively (automatically) adjusted throughout training (here we simply picked reasonable defaults based on Fig. S3 in Hess et al. [2023]). Hess et al. [2023] prove that an optimal choice for $\alpha$ leads to bounded gradients as sequence length $T$ goes to infinity. An extension of GTF to multimodal and non-Gaussian data (multimodal teacher forcing, MTF) has recently been proposed [Brenner et al., 2024]. MTF can be naturally incorporated into our hierarchical framework and hence is an interesting avenue for further applications.
A.1.2 Differential learning rates
For training we used the RAdam [Liu et al., 2020] optimizer. To ensure successful training, we found it crucial to assign a significantly higher learning rate to the subject-specific feature vectors ($10^{-3}$) than to the group-level matrices ($10^{-4}$). This helped prevent numerical instabilities in the group-level matrices, which occurred with higher learning rates. It also prioritizes the incorporation of subject-specific information through the feature vectors, see Fig. 19.
A.1.3 Initialization
We used Xavier uniform initialization [Glorot and Bengio, 2010] for the group-level matrices, which is designed to keep the variance of the outputs $n_{out}$ of a layer roughly equal to the variance of its inputs $n_{in}$ by drawing weights from a uniform distribution in the interval $[-a,a]$, where $a=\sqrt{\frac{6}{n_{in}+n_{out}}}$. We reduced the weights further by a factor of $0.1$ to stabilize training. Additionally, we applied $L_{2}$ regularization to the group-level matrices to further enhance stability.
A.1.4 Batching
We employed a batching strategy that subsamples training sequences $\bm{x}_{\tau:\tau+T_{\text{seq}}^{(j)}-1}$ of length $T_{\text{seq}}^{(j)}$ from the total time series of length $T_{\text{max}}^{(j)}$, starting at a randomly drawn position $\tau\in\{1,T_{\text{max}}^{(j)}-T_{\text{seq}}^{(j)}+1\}$.
@@ A.1.4 Batching
Given the complexity of the datasets considered here, an optimal data loading strategy for stochastic optimization is not obvious a priori. We hence investigated how different batch sizes impact the reconstruction performance of models trained on the 64-subject Lorenz-63 dataset (see main text). The blue curve in Fig. 9 illustrates how performance varies across a range of batch sizes when training with 50 batches per epoch, corresponding to a fixed number of gradient updates per epoch. In this setup, models trained with smaller batch sizes naturally see less data per epoch than those with larger batches. To ensure that the observed effects are not merely due to reduced data exposure with smaller batch sizes, we repeated the experiment using the same batch sizes, but ensuring the entire dataset was seen in each epoch (orange curve). Our findings suggest that smaller batch sizes do not contain enough statistical information across all subjects, leading to higher variance in gradient updates and, in some cases, failure of training. For larger batch sizes, the performance difference was marginal.
@@
The previous experiments were conducted with a fixed number of subjects. However, since we aim to incorporate information from all subjects within each batch, we extended the analysis to examine how the optimal batch size scales with different numbers of subjects. To do this, we randomly sampled smaller subsets from the dataset and repeated the experiment. Fig. 10 gives a heat map of the results, showing that the optimal batch size falls between 4 to 16 times the number of subjects in the dataset. For datasets with a large number of subjects, this range becomes computationally infeasible, so we instead sample batch sequences from a random subset of subjects on each iteration.
A.1.5 Robustness
Figure 11: a: Cross-correlation matrix of cosine similarities between the $S=20$ feature vectors for ten Lorenz-96 training runs. b: Mean performance measures across all $20$ subjects for ten Lorenz-96 training runs.
As evident from Tab. 1, repeated training runs with the hier-shPLRNN deliver highly consistent results. Fig. 11b dissects this further by plotting the median performance (across the $S$ subjects) for all ten training runs separately, which exhibits only little variation. We further quantified the robustness of the learned feature spaces by first calculating the pairwise cosine similarities between subject feature vectors in any given run, and then the correlations among those across ten different training runs (yielding the cross-correlation matrix in Fig. 11a for the Lorenz-96 benchmark). Average cross-correlations for all the different settings studied here are in Tab. 2.
Table 2: Mean $\pm$ SD of feature space correlations for all experiments.
Experiment Lorenz-63 Lorenz-96 Pulse Wave EEG fMRI
Correlation $0.90\pm 0.03$ $0.75\pm 0.05$ $0.88\pm 0.04$ $0.81\pm 0.03$ $0.85\pm 0.05$
A.1.6 Other architectures tested
Besides the shPLRNN for which results are reported in the main text, we repeated the Lorenz-63 experiment in sect. 4.1 with the following RNN architectures trained by GTF [Hess et al., 2023]: First, a simple ‘vanilla RNN’,
$\bm{z}_{t}=\tanh(\bm{W}\bm{z}_{t-1}+\bm{b}),$ (7)
and, second, an LSTM
$\begin{split}\bm{f}_{t}&=\sigma\left(\bm{W}_{f}\bm{x}_{t-1}+\bm{U}_{f}\bm{z}_{t-1}+\bm{b}_{f}\right)\\ \bm{i}_{t}&=\sigma\left(\bm{W}_{i}\bm{x}_{t-1}+\bm{U}_{i}\bm{z}_{t-1}+\bm{b}_{i}\right)\\ \bm{o}_{t}&=\sigma\left(\bm{W}_{o}\bm{x}_{t-1}+\bm{U}_{o}\bm{z}_{t-1}+\bm{b}_{o}\right)\\ \bm{\tilde{c}}_{t}&=\tanh\left(\bm{W}_{f}\bm{x}_{t-1}+\bm{U}_{f}\bm{z}_{t-1}+\bm{b}_{f}\right)\\ \bm{c}_{t}&=\bm{f}_{t}\odot c_{t-1}+\bm{i}\odot\bm{\tilde{c}}_{t}\\ \bm{z}_{t}&=\bm{o}_{t}\odot\tanh\left(\bm{c}_{t}\right)\end{split}$ (8)
where $\odot$ denotes element-wise multiplication. Since both architectures use sigmoidal activation functions, we employ a linear observation model, shared across subjects, to scale the latent states back to the observation range.
Results in Tab. 3 indicate that while both these RNNs also profit from the hierarchisation, performance for the LSTM is considerably worse than for the shPLRNN and vanilla RNN (potentially because due to its specialized architecture the LSTM gains less from the GTF-based training; for DSR, the training technique is actually more crucial than the architecture, Hess et al. [2023]).
Additionally, we tested how using a linear observation model, i.e. $\hat{\bm{x}}_{t}^{(j)}=h_{\bm{\theta}_{\text{obs}}^{(j)}}(\bm{z}_{t}^{(j)})=\bm{B}^{(j)}\bm{z}_{t}^{(j)}$, influences hier-shPLRNN performance. We tested both, using a shared observation matrix across all subjects ($\bm{B}^{(j)}=\bm{B}\ \forall j$), as well as subject-specific observation matrices constructed from the respective feature vectors ($\bm{B}^{(j)}=\text{mat}(\bm{l}^{(j)}\cdot\bm{P}_{B},N,M)$). Results in Tab 3 indicate that both variants of a linear observation model degrade performance slightly.
Table 3: Median ($\pm$ MAD) reconstruction performance for the two alternative RNN architectures, and for two variants of linear observation models for the shPLRNN, on the Lorenz-63 benchmark from Sect. 4.1.
Architecture $D_{\text{stsp}}$ $D_{H}$
Vanilla ensemble $1.72\pm 0.24$ $0.296\pm 0.008$
hierarchical $0.57\pm 0.09$ $0.199\pm 0.010$
LSTM ensemble $2.7\pm 0.3$ $0.76\pm 0.017$
hierarchical $2.3\pm 0.3$ $0.37\pm 0.06$
hier-shPLRNN $0.55\pm 0.26$ $0.087\pm 0.012$
+ shared obs. model
hier-shPLRNN $0.54\pm 0.18$ $0.087\pm 0.010$
+ hierarchical obs. model
A.1.7 Hierarchization scheme
@@
As discussed in Sect. 3, our hierarchization framework is flexible, allowing for different forms of the group-level function $G_{\bm{\theta}_{\text{group}}}$ that maps subject-specific feature vectors to model weights. Here we discuss different choices for parameterizing $G_{\bm{\theta}_{\text{group}}}$ and points to consider in this context. The naive approach, employing direct linear projections, results in over-parameterization. For example, the weight matrices $\bm{W}_{1}$ and $\bm{W}_{2}$ of the shPLRNN have $M\times L$ degrees of freedom, but when parameterized with feature vectors through $\bm{W}_{1}^{(j)}:=\text{mat}(\bm{l}^{(j)}\cdot\bm{P}_{W_{1}},M,L)$, the number of trainable parameters expands to $N_{\text{feat}}\times(M\times L+1)$, which with $N_{\text{feat}}>1$ exceeds the $M\times L$ available entries in $\bm{W}_{1}$.
@@
(analogously for $\bm{W}_{2}$), while the diagonal of $\bm{A}$ and vectors $\bm{h}_{1},\bm{h}_{2}$ were specified directly through linear maps. This reduces the number of parameters to $N_{\text{feat}}\times(M+L+1)$. In the cases where $N_{\text{feat}}<M$, such that the matrices would not be full rank, we applied an extra step to map the feature vector to an auxiliary vector $\mathbb{R}^{M}\ni\bm{a}^{(j)}=\bm{R}\bm{l}^{(j)}$ via the shared [^4] To avoid increasing the number of free parameters, $\bm{R}$ could also be chosen to be random and constant. matrix $\bm{R}\in\mathbb{R}^{M\times N_{\text{feat}}}$, and subsequently mapped to the weights $\bm{W}_{1}^{(j)}=\bm{P}_{W_{1}}\text{diag}(\bm{a}^{(j)})\bm{Q}_{W_{1}}$ (analogously for $\bm{W}_{2}$). We also tested using an MLP for mapping the feature vectors onto the subject-specific DSR model weights, with MLP parameters shared among subjects:
@@
Outer product $3348$ $0.36\pm 0.03$ $0.99$ $0.84$ $0.0986\pm 0.0025$ $-0.63$ $0.27$
MLP $36713$ $0.85\pm 0.09$ $-7.55$ $0.0$ $0.111\pm 0.004$ $-5.54$ $0.0$
A.1.8 Performance as a function of data size
@@
Figure 12: Relative performance (hierarchical/baseline) as a function of time series length for the 64-subject Lorenz-63 dataset (lower = better).
To further evaluate how performance scales with the number of subjects and available time series length, we conducted additional experiments on the Lorenz-63 dataset as in Sect. 4.1, but with a varying number of time steps for each subject/system, where $1000$ training time steps corresponds to the setting in Sect. 4.1. Fig. 12 shows the ratio between the hierarchical model’s performance and the baseline performance (i.e., training an ensemble of individual shPLRNNs, with values below $1$ indicating superior performance of the hierarchical model). Even for long training sequences the hierarchical model consistently outperforms the baseline, particularly in terms of state-space divergence.
We furthermore evaluate how the performance depends on the number of subjects $S$ for different values of time series length $T_{\text{max}}$, using at least two different dynamical regimes (limit cycle and chaos) for the Lorenz-63 (hence starting at $S=2$). Results are in Fig. 13, illustrating that performance gains are substantial up to $S=16$ for this dataset.
Figure 13: Dependence of model performance on the number of subjects $S$ and trajectory length $T_{\text{max}}$.
A.1.9 Computational complexity
Qua construction, inference times (i.e., forward-iterating the full model from some initial condition) should scale linearly with number of subjects $S$ and number of features $N_{\text{feat}}$, which Fig. 14b-c confirms they do (note that due to construction of the subject-specific model matrices $\bm{\theta}^{(j)}_{\text{DSR}}$, initially setup times for the hier-shPLRNN are slightly larger than for an ensemble of shPLRNNs, but this difference is negligible in practice for realistic settings with $N_{\text{feat}}<<S$ and more than made up for by the fast split-second fine-tuning of the hier-shPLRNN once trained).
More importantly, as shown in Fig. 14a, actual training times for the hier-shPLRNN until a pre-set performance criterion was reached ($D_{\text{stsp}}\leq 1.0$), actually decrease with the number of subjects $S$ (note that all training here was intentionally performed on a single CPU core, hence there was no parallelization here). The reason for this presumably is that incorporation of more subjects, and hence sampling of a larger portion of the underlying systems’ parameter space, allows the model to reach the criterion and reconstruct each subject’s dynamics much more quickly. This further illustrates that the hier-shPLRNN leverages and transfers information from multiple subjects highly efficiently.
Figure 14: a: Wall clock time of model convergence (i.e. first passage time for median $D_{\text{stsp}}\leq 1$) for models trained with different $S$ (using the Lorenz-63 system with $\rho\in\{28,34,40,46,52,58,64,70,76,82\}$). For all values of $S$, the total number of data points was held constant at $20,000$ time steps, uniformly divided among all subjects. Error bars = SEM across ten runs, sampling $D_{\text{stsp}}$ every 50 epochs. All models were trained on a single core of an Intel Xeon Gold 6254. b: Wall clock time of model inference ($T_{\text{seq}}=30$, $M=3$, $L=50$) for different $N_{\text{feat}}$ at $S=2048$. c: Wall clock time of model inference for different $S$ and $N_{\text{feat}}$.
A.1.10 Uncertainty Estimates
@@
To obtain the uncertainty estimates (error bars) in Figs.6b and c, we first drew a long trajectory with $10000$ time steps from the ground truth system using ${\rho}^{(S+1)}$. We then randomly sampled multiple training trajectories of length $T_{\text{max}}$, inferred the feature vectors $\hat{\bm{l}}^{(S+1)}$ for each of those individually, and calculated the standard deviation of the resulting $\hat{\rho}^{(S+1)}$ estimates. As shown in Fig. 15, the estimates sharpen as longer training sequences are used.
@@
Figure 15: a: Distributions $p(\hat{\rho}^{(j)}\mid\bm{x}_{1...T_{\text{max}}}^{(j)})$ (approximated using kernel density estimates), estimated from $10$ time series each, with different lengths $T_{\text{max}}$ (color-coded) for the example $\rho^{(j)}_{\text{test}}$ from Fig. 6. b: Corresponding GTF loss curves for different training sequence lengths, illustrating that longer sequences/more training data lead to smaller standard deviation over loss curves.
A.1.11 Hyperparameters